OpenStudy (anonymous):

could anyone give an example of solving an equation with square root expressions on both sides of the equation

jimthompson5910 (jim_thompson5910):

Something like this? \[\Large \sqrt{2x+5} = \sqrt{3x-1}\]

OpenStudy (anonymous):

I'll try that, would you know what a real world example would be?

jimthompson5910 (jim_thompson5910):

here is a page that has some applications of radicals (ie square roots) http://www.regentsprep.org/Regents/math/algtrig/ATE10/radModelprac.htm